OpenStudy (anonymous):

Write 3cos(x) + 2isin(x) in a form involving the (x+iy) factor... Help?! Thanks!

OpenStudy (anonymous):

Your formula sent me on a wild goose chase for a while, because I interpreted sin(2x) / cosh(2y)+cos(2x) as: sin(2x) ------- + cos(2x) cosh(2y) instead of: sin(2x) ----------------- cosh(2y)+ cos(2x) I would expect to see this last expression written as sin(2x)/(cosh(2y)+cos(2x)) Ahlfors' argument, as he admits, is heuristic, and is more of a mnemonic, but he assures the reader that it can be proved. In this case, u(0,0) = 0, and the key is that for an imaginary argument iw (w real), e^(iw) + e^(-iw) cosh(iw) = ---------------- 2 (you get this from the formula for cosh(x) when x is real, by just replacing x by a complex value, in this case, a pure imaginary one). Using Euler's formula: e^(it) = cos(t) + i*sin(t) cos(w) + i*sin(w) + cos(w) - i*sin(t) cosh(iw) = ------------------------------------- 2 = cos(w) Also, cosh(-iw) = cosh(i*(-w)) = cos(-w) = cos(w) Since i^4 = 1 1 - = -i i so f(z) = 2*u(z/2,z/(2i)) - u(0,0) sin(2*(z/2)) = 2 * -------------------------- - 0 cosh(2*(z/2i))+cos(2*(z/2) sin(z) = 2 * ------------------ cosh(-iz) + cos(z) sin(z) = 2 * --------------- cos(z) + cos(z) = tan(z)

OpenStudy (anonymous):

w=3cos(x) + 2isin(x)= Ln|w| +iArg(w)

OpenStudy (anonymous):

Woah Rohangrr! Impressive reply, but you lost me from the get go! .. How did you get sin(2x) / cosh(2y)+cos(2x) in the first place?

OpenStudy (anonymous):

@feldy90 it was just to make easy for u!

OpenStudy (anonymous):

nothing else!

OpenStudy (anonymous):

see the work done

OpenStudy (anonymous):

okay.. cheers!

OpenStudy (anonymous):

feldy u must have give reward ft that

OpenStudy (anonymous):

i did...

OpenStudy (anonymous):

thanx!